A chemist needs a 20% solution of alcohol. How many liters of a 10% alcohol solution must be mixed with 90 liters of a 40% solution to get a 20% solution?
Answer by CharlesG2(834)

A chemist needs a 20% solution of alcohol. How many liters of a 10% alcohol solution must be mixed with 90 liters of a 40% solution to get a 20% solution?
% in A * amount A + % in B * amount B = % in C * amount C
0.10 * x + 0.40 * 90 = 0.20 * (x + 90)
0.10x + 36 = 0.20x + 18
-0.10x = -18
0.10x = 18
x = 180 liters